### v2

# InceptionMamba: An Efficient Hybrid Network with Large Band Convolution and Bottleneck Mamba

Abstract
Within the family of convolutional neural networks, InceptionNeXt has shown excellent competitiveness in image classification and a number of downstream tasks. Built on parallel one-dimensional strip convolutions, however, it suffers from limited ability of capturing spatial dependencies along different dimensions and fails to fully explore spatial modeling in local neighborhood. Besides, inherent locality constraints of convolution operations are detrimental to effective global context modeling. To overcome these limitations, we propose a novel backbone architecture termed InceptionMamba in this study. More specifically, the traditional one-dimensional strip convolutions are replaced by orthogonal band convolutions in our InceptionMamba to achieve cohesive spatial modeling. Furthermore, global contextual modeling can be achieved via a bottleneck Mamba module, facilitating enhanced cross-channel information fusion and enlarged receptive field. Extensive evaluations on classification and various downstream tasks demonstrate that the proposed InceptionMamba achieves state-of-the-art performance with superior parameter and computational efficiency. The source code will be available at https://github.com/Wake1021/InceptionMamba.
journal: Pattern Recognition
(a) Params. vs Top-1(%)
(b) FLOPs vs Top-1(%)
Figure 1: Performance on ImageNet-1K validation set at $224^{2}$ resolutions.
1 Introduction
In Convolutional Neural Networks (CNNs) [1, 2, 3], the success of Inception architecture is largely attributed to efficient strip convolution, evolving from local feature extraction to dynamic parameter optimization [3]. Originated from InceptionNeXt [4], it is beneficial for reducing the number of parameters through asymmetric convolution decomposition (such as splitting a 7×7 convolution into 1×7 and 7×1 strip convolutions). As a representative member within Inception family, InceptionNeXt has showcased the efficacy of strip convolutions in balancing computational efficiency and representational abilities. In InceptionNeXt, large-kernel depthwise convolutions are decomposed into four efficient branches including parallel one-dimensional strip convolutions, such that large-kernel CNNs are accelerated with maintained performance.
Despite achieving huge success, conventional one-dimensional strip convolutions in InceptionNeXt suffer twofold inherent limitations, which hinder their potential in complex visual understanding scenarios. On the one hand, the one-dimensional strip convolutions are independently computed in separate channel groups, lacking sufficient spatial modeling capacity in capturing complex visual contents. For instance, while horizontal kernels effectively characterize axis-aligned patterns, their weakened responses to orthogonal structures (e.g., vertical edges) results in less discriminative encoding. Additionally, the local inductive bias of CNNs, arising from utilizing limited-size convolutional kernels (e.g., 3$\times$3 convolutions) for hierarchical local feature extraction, fundamentally constrains their capacity in global contextual relationship modeling, thereby suffering from restricted receptive field. While recent efforts to addressing this limitation have explored oversized kernels [5] (e.g., 31$\times$31 convolutions) and dilated convolutions [6] to expand receptive fields, these approaches are challenged by accuracy-efficiency trade-off: enlarged kernels introduce a quadratic increase in computational complexity $(O(k^{2}))$, while dilated convolutions are prone to grid artifacts that degrade spatial coherence in feature representations.
Figure 2: Visualizations of different CAMs [7] generated from ConvNeXt(b), InceptionNeXt(c) and our InceptionMamba(d) given a panda image(a). It can be observed that InceptionMamba showcases superior spatial modeling capacity in capturing semantic-aware areas in the neighborhood of panda eye and body. In contrast, InceptionNeXt only focuses on scattered eye-centric regions, yielding less cohesive representations.
To overcome the aforementioned drawbacks, we propose an efficient hybrid backbone architecture with large band convolutions and bottleneck Mamba termed InceptionMamba. Inherited from Inception architecture, our InceptionMamba maintains Inception-style depthwise convolutions with efficient parallel structures, yet features carefully designed orthogonal band convolutions instead of the traditional one-dimensional stripe convolutions. Thus, multi-dimensional visual cues can be discovered for cohesive spatial modeling, leading to comprehensive embedding with improved representational capacity. In addition, to improve inter-channel interaction ability, we incorporate a bottleneck Mamba module into the token mixture for facilitating cross-channel information integration and enlarged receptive field. The advantages of our framework are illustrated in Fig. 2, demonstrating that InceptionMamba can capture semantically correlated local neighborhood with sufficient discriminative power. For instance, our method can accurately focus on object-aware regions surrounding panda eye and body, whereas InceptionNeXt only identify isolated eye-centric areas with degraded spatial modeling capacity.
The superiority of our InceptionMamba has been verified through extensive evaluations on various public benchmarking datasets for both classification and other downstream tasks. As shown in Fig. 1, our InceptionMamba-B attains 84.7% top-1 accuracy on ImageNet-1K, outperforming state-of-the-arts including both InceptionNeXt and Mamba-like frameworks (e.g., MambaVision [8] and VMamba [9]), while enjoying competitive computational efficiency compared to InceptionNeXt.
To summarize, our contributions in this study are fourfold as follows:
  • 1.
    We propose a hybrid backbone architecture termed InceptionMamba with efficient parallel structure inherited from Inception and develop a model family accordingly.
  • 2.
    In contrast to one-dimensional strip convolution involved in InceptionNeXt [4], our carefully designed orthogonal band convolutions significantly contribute to characterizing semantically cohesive visual cues in local adjacent neighborhood, thereby demonstrating improved spatial modeling capacity.
  • 3.
    The devised bottleneck Mamba module is conducive to inter-channel information fusion, while expanding the receptive field for enhancing global contextual modeling.
  • 4.
    Extensive evaluations on public benchmarking datasets for classification and other downstream tasks demonstrate that our InceptionMamba achieves state-of-the-art performance with competitive efficiency.
2 Related Work
2.1 Inception Architecture
The evolution of the Inception architecture began with the core idea of multi-scale feature fusion. Earliest Inception architecture [10] introduces parallel multi-branch convolutions (1$\times$1, 3$\times$3, 5$\times$5, and pooling) along with 1$\times$1 dimensionality reduction to overcome the single-scale limitation. Subsequent improvements of Inception-v2 [11] and Inception-v3 [11] are manifested in enhancing computational efficiency via convolution decomposition (e.g., splitting one 5$\times$5 convolution into two 3×3 convolutions) and batch normalization (BN). Similar pure Inception variant is Inception-v4 [12] which contains multiple consecutive Inception blocks with parallel structures. More advanced Inception-ResNet is a hybrid Inception network which integrates the residual learning into Inception architecture, achieving network training speedup through skip connections. A milestone work within Inception family emerges as InceptionNeXt [4] which decomposes large-kernel depthwise convolutions into multiple parallel branches of smaller kernels (e.g., square 3$\times$3 convolutions and one-dimensional strip convolutions), achieving desirable accuracy-throughputs trade-off in a series of downstream tasks. It is significantly beneficial for accelerating large-kernel networks with maintained performance.
2.2 Mamba
As a sequence modeling framework inspired by control theory, Mamba [13, 14] is a new State Space Model (SSM) which is capable of capturing long-range dependencies through the evolution of hidden states. The continuous system is defined by three core matrices: the state transition matrix $\mathbf{A}\in\mathbb{R}^{M\times M}$, the input projection matrix $\mathbf{B}\in\mathbb{R}^{M\times 1}$, and the output projection matrix $\mathbf{C}\in\mathbb{R}^{1\times M}$. The system dynamics can be described as follows:
$h^{{}^{\prime}}(t)$ $=\mathbf{A}h(t)+\mathbf{B}x(t)$ (1)
$y(t)$ $=\mathbf{C}h(t)$
To achieve efficient discrete computation, Mamba [13, 14] introduces zero-order preserving discretization method, which converts the continuous system into a recursive form through the time scale parameter $\Delta$ :
$\mathbf{\overline{A}}$ $=e^{\Delta\mathbf{A}}$ (2)
$\quad\mathbf{\overline{B}}$ $=({\Delta\mathbf{A}})^{-1}(e^{{\Delta\mathbf{A}}}-\mathbf{I})\cdot{\Delta\mathbf{B}}$
Thus, the discrete state update equation is obtained:
$h_{t}$ $=\mathbf{\overline{A}}h_{t-1}+\mathbf{\overline{B}}x_{t}$ (3)
$\quad y_{t}$ $=\mathbf{C}h_{t}$
Retaining the sequence modeling characteristics of RNN, the above formulation can be reconstructed in a parallel computing manner from the global convolution perspective. The system output can be equivalently expressed as:
$\mathbf{\overline{K}}$ $=(\mathbf{C\overline{B}},\mathbf{C\overline{AB}},\dots,\mathbf{C\overline{A}^{L-1}\overline{B}})$ (4)
$\mathbf{y}$ $=\mathbf{x}\ast\overline{\mathbf{K}}$
where $\ast$ represents the convolution operation, and $\overline{\mathbf{K}}\in\mathbb{R}^{L}$ serves as the kernel of the SSM.
Recent advancements in visual Mamba models demonstrate the potential of SSM in a variety of vision tasks. In Vision Mamba [15], a bidirectional Mamba block was proposed for efficient high-resolution image processing. VMamba [9] introduces Cross-Scan strategy to strengthen global context modeling. MambaVision [8] combines linear complexity of Mamba with Transformer self-attention in a hybrid architecture to enhance efficiency and global spatial reasoning. Recent research explores the essence of Mamba, and reveals that Mamba is not ideally suited for all vision tasks. Meanwhile, a range of MambaOut [16] models are obtained by stacking Mamba blocks while removing their core token mixer SSM.
Figure 3: Our InceptionMamba architecture with four stages. Similar to ConvNeXt [3], InceptionMamba employs hierarchical architecture of four consecutive stages. Each stage consists of a patch embedding layer or a downsampling module, combined with $N_{i}$ InceptionMamba blocks.
3 Method
Encouraged by recent success of vision Mamba models, we propose to deeply integrate Mamba SSM [13, 14] into efficient InceptionNeXt [4] architecture with parallel structures. The former, which enjoys linear-complexity sequence modeling capability, can capture long-range spatial dependencies, whilst the latter excels at extracting key features while maintaining network efficiency. In this sense, our synergistic design helps to preserve strengths of CNNs in efficient local reasoning while enhancing model adaptability to complex scenes through global dynamic feature selection mechanism of Mamba [13, 14], establishing a new foundation paradigm for a series of downstream vision tasks.
3.1 Overview of InceptionMamba
Inspired by InceptionNeXt [4], we have designed a four-stage InceptionMamba architecture as illustrated in Fig. 3. For a given input tensor $X\in\mathbb{R}^{B\times C\times H\times W}$ where $B$, $C$, $H$, $W$ represent the batch size, number of channels, height, and width, respectively, the feature processing begins with a Local Mixer module that emphasizes local spatial information encoding similar to token mixer in InceptionNeXt [4]:
$X^{{}^{\prime}}=\text{LocalMixer}(X)$ (5)
The output $X^{{}^{\prime}}$ is then forwarded to a Global Mixer module for global context modeling and capturing long-range dependencies:
$X^{{}^{\prime\prime}}=\text{GlobalMixer}(X^{{}^{\prime}})$ (6)
Then, normalization is applied to the aggregated feature expressed as:
$Y=\text{Norm}(X^{{}^{\prime\prime}})$ (7)
After normalization, the resulting feature $Y$ is passed through a MLP module consisting of two fully connected layers, with a non-linear activation function inserted in between. These fully connected layers are typically implemented as $1\times 1$ convolutions for efficiency. Additionally, a residual connection is employed to facilitate gradient flow and model stability. This process is formulated as:
$Y$ $=\text{Conv}_{1\times 1}^{rC\to C}\{\sigma[\text{Conv}_{1\times 1}^{C\to rC}(Y)]\}+X$ (8)
where $\text{Conv}_{k\times k}^{C_{i}\to C_{o}}$ denotes a convolution operation with kernel size $k\times k$. Input and output channels are denoted as $C_{i}$ and $C_{o}$, respectively. Additionally, $r$ is the expansion ratio and $\sigma$ denotes a non-linear activation function.
As the cornerstone of our method, the core InceptionMamba block constitute four-stage framework resembling ConvNeXt [3] and InceptionNeXt [4], leading to a family of InceptionMamba model with varying sizes. More specifically, the block numbers in the three model variants are [3, 3, 12, 3], [4, 4, 32, 4] and [4, 4, 34, 4] for the Tiny (T), Small (S) and Base (B), respectively. Starting from an input image of $3\times H\times W$, our model first applies patch embedding to process the visual data. In each stage, subsequently, downsampling operation is combined with multiple InceptionMamba blocks to progressively reduce spatial resolution, i.e., spatial size $H\times W$ is reduced by 4 times, 16 times, 32 times, and 64 times respectively, while expanding the channel dimension (C). This pyramid-like hierarchical architecture ends with global average pooling and classifiers for the final prediction. The architecture configurations of the three InceptionMamba variants can be referred to Table 1.
Table 1: Architectures of our InceptionMamba family including Tiny (T), Small (S) and Base (B) versions.
Stage #Tokens Layer Specification InceptionMamba
T S B
1 $\frac{H}{4}\times\frac{W}{4}$ PatchEmbed Kernel Size $3\times 3$, stride 2
Embed. Dim. 72 96
InceptionMamba Block Kernel size $3\times 3,3\times 11,11\times 3$
Conv. group ratio 1/8
Bottleneck ratio 1/2
MLP Ratio 4
# Block 3 4
2 $\frac{H}{8}\times\frac{W}{8}$ DownSampling Kernel Size $3\times 3$, stride 2
Embed. Dim. 144 192
InceptionMamba Block Kernel size $3\times 3,3\times 11,11\times 3$
Conv. group ratio 1/8
Bottleneck ratio 1/2
MLP Ratio 4
# Block 3 4
3 $\frac{H}{16}\times\frac{W}{16}$ DownSampling Kernel Size $3\times 3$, stride 2
Embed. Dim. 288 384
InceptionMamba Block Kernel size $3\times 3,3\times 11,11\times 3$
Conv. group ratio 1/8
Bottleneck ratio 1/2
MLP Ratio 4
# Block 12 32 34
4 $\frac{H}{32}\times\frac{W}{32}$ DownSampling Kernel Size $3\times 3$, stride 2
Embed. Dim. 576 768
InceptionMamba Block Kernel size $3\times 3,3\times 11,11\times 3$
Conv. group ratio 1/8
Bottleneck ratio 1/2
MLP Ratio 4
# Block 3 4
Global average pooling, MLP
Parameters (M) 25 46 83
FLOPs (G) 4.0 7.8 14.3
Figure 4: Illustrative comparison of ConvNeXt [3], InceptionNeXt [4], and our InceptionMamba block. InceptionMamba consists of two key components, namely LocalMixer and GlobalMixer. The former inherits efficient parallel structure from InceptionNeXt [4] to accelerate large-kernel depthwise convolutions. Different from InceptionNeXt [4] using one-dimensional strip convolutions, our InceptionMamba takes advantage of orthogonal band convolutions to capture multi-dimensional visual cue for cohesive feature embedding. Furthermore, the GlobalMixer adopts a carefully designed bottleneck Mamba [13, 14] to facilitate inter-channel information with enlarged receptive field. Compared to ConvNeXt [3] and InceptionNeXt [4], our InceptionMamba enjoys efficient local spatial modeling capacity while excels in capturing long-range dependencies.
3.2 LocalMixer with large-kernel band convolutions
Inspired by SLaK [17] and InceptionNeXt [4], we devise a multi-branch structure in our LocalMixer to address the limitation of one-way large kernels. Inherited from InceptionNeXt[4] block, the parallel structure allows speedup of large-kernel depthwise convolutions for efficient local spatial modeling. Different from its predecessor, however, our LocalMixer employs orthogonal large-kernel band convolutions instead of one-dimensional strip convolutions in InceptionNeXt [4]. Band convolution can focus on larger areas with improved local modeling capability. This advantage can be intuitively shown in Fig. 2, clearly demonstrating that our InceptionMamba generates more spatially cohesive activations on the object-aware regions compared to InceptionNeXt[4].
Mathematically, the input features $X$ are divided into three groups along the channel dimension via our LocalMixer as follows:
$X_{square},X_{band},X_{identity}=Split(X)$ (9)
which are subsequently fed into three different branches for parallel processing:
$X_{1}$ $=DWConv_{3\times 3}(X_{square})$ (10)
$X_{2}$ $=DWConv_{3\times 11}(X_{band})+DWConv_{11\times 3}(X_{band})$
$X_{3}$ $=X_{id}$
In Eq. (10), the sizes of orthogonal large-kernel band convolutions are $3\times 11$ and $11\times 3$. Finally, the output features of each branch are concatenated as:
$X^{{}^{\prime}}=Concat(X_{1},X_{2},X_{3})$ (11)
Notably, identity mapping operations are applied to the majority of channel groups in the LocalMixer while depthwise convolutions are only imposed on a small subset of groups. This is consist with InceptionNeXt [4] which reveals that depthwise convolutions are not required for most channels.
3.3 GlobalMixer with bottleneck Mamba
After the LocalMixer, we design a GlobalMixer module for inter-channel interaction at the Token layer. Traditional methods including ConvNeXt [3] and InceptionNeXt [4] only pass the concatenated features through a straightforward MLP, leading to inadequate cross-channel interactions due to inherent single-channel processing characteristic of depthwise convolution. To mitigates this drawback, a bottleneck structure is introduced into the GlobalMixer, making use of $1\times 1$ convolution to compress and expand feature channels. This is beneficial for efficient cross-channel fusion in a low-dimensional space while retaining the key information of the original features.
To strengthen global modeling capacities while optimizing computational efficiency, a state-space module (SS2D) [9] is integrated within the bottleneck architecture. This co-design leverages channel compression mechanism (ratio $r=2$) of the bottleneck for significantly reducing the SS2D [9] dimensionality. More importantly, the SS2D [9] module can capture long-range dependencies with the help cross-scan mechanism, while the compressed feature space enjoys desirable parameter efficiency. Mathematically, given the Input $X^{{}^{\prime}}$, our GlobalMixer with a bottleneck Mamba [13, 14] is formulated as follows:
$X^{{}^{\prime\prime}}$ $=\text{Conv}_{1\times 1}^{C\to C/r}(X^{{}^{\prime}}))$ (12)
[…]
$X^{{}^{\prime\prime}}$ $=\text{Conv}_{1\times 1}^{C/r\to C}(X^{{}^{\prime\prime}})$
$Y$ $=X+X^{{}^{\prime}}$
[…]
3.4 Block Comparisons
Since our InceptionMamba is closely related to ConvNeXt [3] and InceptionNeXt [4] frameworks, we compare different building blocks comprising the three frameworks as illustrated in Fig. 4. With the help of the LocalMixer, our InceptionMamba block allows efficient computation of large-kernel depthwise convolutions via a parallel multi-branch structure inherited from the InceptionNeXt [4]. Different from the predecessor which makes use of one-dimensional strip convolutions, our method features orthogonal band convolutions with improved local modeling capacity in characterizing multi-dimensional visual cues. Additionally, the GlobalMixer module with bottleneck Mamba [13, 14] structure can capture long-range dependencies through cross-channel interaction, which is conducive to boosting global modeling capability with expanded receptive field. In contrast, both ConvNeXt [3] and InceptionNeXt [4] demonstrate limited global modeling capacity, lacking GlobalMixer-like component for sufficient information fusion.
4 Experiments
To evaluate the proposed InceptionMamba architecture, we have carried our extensive experiments for classification and other downstream vision tasks including object detection and semantic segmentation.
4.1 Image classification on ImageNet
Experimental setup
For classification, the public ImageNet-1k [18, 19] dataset is involved in our evaluations. Our model is trained from scratch at 224$\times$224 resolution for 300 epochs using the AdamW optimizer with a weight decay of 5e-2, initial learning rate of 1e-3, and batch size of 512. The training strategy incorporates a cosine scheduler with 10 warmup epochs, Label Smoothing of 0.1, stochastic depth, and RandAugment. For fairness, we adopt conventional data augmentation methods including Mixup, random erasing, and AutoAugment following [20]. Numerous advanced state-of-the-art backbone networks are involved in our comparative studies, including Vision-RWKV [21], Vision Mamba [15], Vision LSTM [22], PlainMamba [23], EfficientVMamba [24], VMamba [9], InceptionNeXt [4], MambaVision [8], QuardMamba [25], PeLK [26], MambaOut [16] and RDNet [27]. Parallel training is performed with eight NVIDIA RTX 4090 GPUs in all our experiments.
Results
As shown in Table 2, our proposed InceptionMamba demonstrates consistent performance advantages across all model scales while maintaining superior computational efficiency. Among tiny models, InceptionMamba-T achieves 83.1% Top-1 accuracy with only 25.4M parameters and 4.0G FLOPs, outperforming both Mamba and CNN models like VMamba-T (82.2%) and InceptionNeXt-T (82.3%). This trend continues in small and base variants, demonstrating that InceptionMamba-S achieves 84.1% Top-1 accuracy on par with MambaOut-S with lower computational costs (7.8G vs 9.0G FLOPs). Analogously, InceptionMamba-B achieves state-of-the-art performance, reporting unrivaled 84.7% accuracy with only 83M network parameters and 14.3G FLOPs. These results validate the superiority of our InceptionMamba in accuracy-efficiency trade-off compared to existing mainstream backbone networks.
Table 2: Performance comparison on ImageNet-1k for classification across three model scales, i.e., Tiny(T), Small(S) and Base(B).
Model Pub. Reso. Params(M) FLOPs(G) Top-1(%)
VRWKV-S [21] ICLR 2025 224 24 4.6 80.1
Vim-S [15] ICML 2024 224 26 5.1 80.5
ViL-S [22] ICLR 2025 224 23 5.1 81.5
PlainMamba-L2 [23] BMVC 2024 224 25 8.1 81.6
EfficientVMamba-B [24] AAAI 2025 224 33 4.0 81.8
VMamba-T [9] NIPS 2024 224 22 5.6 82.2
InceptionNeXt-T [4] CVPR 2024 224 28 4.2 82.3
MambaVision-T [8] CVPR 2025 224 32 4.4 82.3
QuardMamba-S [25] NIPS 2024 224 31 5.5 82.4
PeLK-T [26] CVPR 2024 224 29 5.6 82.6
MambaVision-T2 [8] CVPR 2025 224 35 5.1 82.7
MambaOut-T [16] CVPR 2025 224 27 4.5 82.7
RDNet-T [27] ECCV 2024 224 24 5.0 82.8
InceptionMamba-T Ours 224 25 4.0 83.1
PlainMamba-L3 [23] BMVC 2024 224 50 14.4 82.3
MambaVision-S [8] CVPR 2025 224 50 7.5 83.3
InceptionNeXt-S[4] CVPR 2024 224 49 8.4 83.5
VMamba-S [9] NIPS 2024 224 50 8.7 83.6
RDNet-S [27] ECCV 2024 224 50 8.7 83.7
QuardMamba-B [25] NIPS 2024 224 50 9.3 83.8
PeLK-S [26] CVPR 2024 224 50 10.7 83.9
MambaOut-S [16] CVPR 2025 224 48 9.0 84.1
InceptionMamba-S Ours 224 46 7.8 84.1
Vim-B [15] ICML 2024 224 98 - 81.9
VRWKV-B [21] ICLR 2025 224 94 18.2 82.0
ViL-B [22] ICLR 2025 224 89 18.6 82.4
VMamba-B [9] NIPS 2024 224 89 15.4 83.9
InceptionNeXt-B [4] CVPR 2024 224 87 14.9 84.0
MambaVision-B [8] CVPR 2025 224 98 15.0 84.2
PeLK-B [26] CVPR 2024 224 89 18.3 84.2
MambaOut-B [16] CVPR 2025 224 85 15.8 84.2
RDNet-B [27] ECCV 2024 224 87 15.4 84.4
InceptionMamba-B Ours 224 83 14.3 84.7
4.2 Object detection &\&& instance segmentation on COCO
Experimental setup
For object detection and instance segmentation, we have conducted experiments on MS-COCO which contains 118k training images and 5k validation images [28]. The detection framework is utilized as Mask R-CNN [29] with our InceptionMamba backbone, and training the detector is initialized with the model pretrained on the ImageNet. In implementation, we follow the standard 1$\times$ training protocol (12 epochs), and adopt multi-scale data augmentation. To maintain the original aspect ratio, the input image is resized such that the shorter side is fixed at 800 pixels while the longer side does not exceed 1333 pixels. Using the AdamW optimizer, initial learning rate is set to 0.0002, whilst weight attenuation coefficient is 0.05. In comparative studies, competing architectures range from CNN and Transformer like ConvNeXt [3] and Swin-Transformer [30] to Mamba-based networks, e.g., EfficientVMamba and MambaOut.
Table 3: Performance comparison on COCO for object detection and instance segmentation. Mask R-CNN is used as detection framework.
Backbone Params(M) FLOPs(G) $AP^{b}$ $AP^{m}$
ConvNeXt-T [3] 48 262 44.2 40.1
Swin-T [30] 48 267 42.7 39.3
ViT-Adapter-S [31] 48 403 44.7 39.9
PVTv2-B2 [32] 45 309 45.3 41.2
EfficientVMamba-B [24] 53 252 43.7 40.2
PlainMamba-L1 [23] 31 388 44.1 39.1
MambaOut-T [16] 43 262 45.1 41.0
InceptionMamba-T 43 233 46.0 41.8
ConvNeXt-S [3] 70 348 45.4 41.8
Swin-S [30] 69 354 44.8 40.9
PVTv2-B3 [32] 65 397 47.0 42.5
MambaOut-S [16] 65 354 47.4 42.7
InceptionMamba-S 63 301 47.5 42.6
ConvNeXt-B [3] 108 486 47.0 42.7
Swin-B [30] 107 496 46.9 42.3
ViT-Adapter-B [31] 102 557 47.0 41.8
PVTv2-B5 [32] 102 557 47.4 42.5
MambaOut-B [16] 100 495 47.4 43.0
InceptionMamba-B 99 421 48.1 43.1
Results
As illustrated in Table 3, our evaluations demonstrate superior performance of InceptionMamba across different model scales on COCO dataset. To be specific, InceptionMamba-T achieves 46.0% $AP^{b}$ and 41.8% $AP^{m}$ with only 43M parameters and 233G FLOPs, which consistently beats other competitors. In particular, InceptionMamba-T surpasses MambaOut-T by almost 1% $AP^{b}$ with superior computational efficiency with 12.5% less FLOPs. This advantage against ConvNeXt-T is elevated to 1.8% $AP^{b}$ with 10.4% less network parameters and 12.5% less FLOPs. At larger model scale, our InceptionMamba-B reports the highest 48.1% $AP^{b}$ and 43.1% $AP^{m}$, which outperforms MambaOut-B and ConvNeXt-B while reducing computational costs by approximately 13% and 15%. In particular, our InceptionMamba achieves better performance than other typical Mamba-based architectures including PlainMamba, EfficientVMamba and MambaOut, sufficiently exhibiting the advantage of our framework in balancing local-aware spatial encoding and global context modeling for capturing long-range dependencies.
Table 4: Comparison of semantic segmentation results on ADE20K dataset using UperNet.
Backbone Params(M) FLOPs(G) mIoU(%)
ConvNeXt-T [3] 60 939 46.0
Swin-T [30] 60 945 44.4
Focal-T [33] 62 998 45.8
MambaVision-T [8] 55 945 46.0
EfficientVMamba-B [24] 65 930 46.5
VMamba-T [9] 55 964 47.3
MambaOut-T [16] 54 938 47.4
InceptionMamba-T 53 928 47.3
ConvNeXt-S [3] 82 1027 48.7
Swin-S [30] 81 1038 47.6
Focal-S [33] 85 1130 48.0
MambaVision-S [8] 84 1135 48.2
VMamba-S [9] 76 1081 49.5
MambaOut-S [16] 76 1032 49.5
InceptionMamba-S 73 1006 49.2
ConvNeXt-B [3] 122 1170 49.1
Swin-B [30] 121 1188 48.1
Focal-B [33] 126 1354 49.0
MambaVision-B [8] 126 1342 49.1
VMamba-B [9] 110 1226 50.0
MambaOut-B [16] 112 1178 49.6
InceptionMamba-B 110 1145 50.1
4.3 Semantic segmentation on ADE20K
Experimental setup
In addition to the above-mentioned evaluations, we perform semantic segmentation experiments on public ADE20K dataset [34] using UperNet architecture [35]. This dataset has 150 semantic categories with 20,000 training images and 2,000 validation images. Initialized with the pretrained model similar to the detection experiments, we follow [16] to adopt a protocol of 160,000 iterations in our training process, and utilize the AdamW optimizer with an initial learning rate set to 0.00012 and a batch size of 16.
Results
Table 4 presents the performance of various architectures on ADE20K dataset. It is shown that our tiny and small InceptionMamba models report respective 47.3% and 49.2% mIoU with superior efficiency, exhibiting competitive performance compared to corresponding VMamba and MambaOut competitors. When extending our framework to larger scales, the resulting InceptionMamba-B achieves 50.1% mIoU with 110M parameters and 1145G FLOPs, which outperforms all the other competing networks in both accuracy and efficiency. These comparisons further reveal the potential of our hybrid architecture in combining local-aware encoding and global context modeling.
4.4 Ablation studies
Branch ratio
To investigate the impact of different branch allocation ratios on model performance, in the structure without the Bottleneck Mamba, we conducted ablation experiments comparing three distinct channel allocation strategies. The experimental results demonstrate that when employing a ratio of [0.125, 0.125, 0.75], the model achieves 81.9% Top-1 accuracy with 21.4M parameters and 3.35G FLOPs computational cost, exhibiting the optimal balance between accuracy and efficiency. In contrast, increasing the ratio of the second branch to 0.25 ([0.125, 0.25, 0.625]) leads to a slight parameter increase to 21.5M and a marginal performance drop (81.7%), indicating that the asymmetric convolutional branches are sensitive to channel variations. A balanced allocation scheme ([0.25, 0.25, 0.5]) maintains 81.9% accuracy with comparable computational cost (3.38G FLOPs), verifying the robustness of the module design to different allocation ratios. After comprehensive evaluation, we select [0.125, 0.125, 0.75] as the final configuration, as it achieves the best trade-off between computational efficiency and model performance.
Table 5: Comparison Results of different Branch Ratio
Branch ratio Params(M) FLOPs(G) Top-1(%)
[0.125, 0.125, 0.75] 21.4 3.35 81.9
[0.125, 0.25, 0.625] 21.5 3.37 81.7
[0.25, 0.25, 0.5] 21.5 3.38 81.9
Local mixer
We conducted comprehensive ablation studies by replacing our token mixer with two established designs: (1) ConvNeXt’s [3] depthwise convolutions and (2) InceptionNeXt’s [4] decomposed convolution. We also replaced band convolution with strip convolution for testing. Experimental results demonstrate that our proposed token mixer achieves superior performance while maintaining identical model complexity.
Table 6: Comparison Results of Different Local Mixers
Local mixer Params(M) FLOPs(G) Top-1(%)
DWConv3$\times$3 25.4 4.0 82.8
InceptionDWConv2d 25.4 4.0 82.9
strip convolution 25.4 4.0 83.0
Ours 25.4 4.0 83.1
Necessity of Bottleneck
The ablation study demonstrates that integrating the Bottleneck structure with SS2D [9] maintains identical model accuracy while significantly improving efficiency. Compared to the baseline SS2D [9], the Bottleneck-enhanced version achieves the same Top-1 performance while reducing both parameters and computational costs by over 10%. Additionally, the throughput increases substantially, highlighting the Bottleneck’s ability to compress feature dimensions effectively without sacrificing performance.
Table 7: Comparsion Results of SS2D and Bottleneck + SS2D
Global mixer Params(M) FLOPs(G) Top-1(%) TP.(imgs/s)
SS2D 29 4.6 83.1 1042
Bottleneck + SS2D 26 4.0 83.1 1362
Module in Bottleneck
The ablation experiments systematically evaluate different bottleneck-integrated architectures, with Bottleneck + SS2D emerging as the optimal configuration. Compared to the baseline without Mamba Bottleneck (21.4M Params, 3.3G FLOPs, 81.9% Top-1), all bottleneck variants improve accuracy, but Bottleneck + SS2D achieves the highest Top-1 accuracy (83.1%) while maintaining competitive efficiency—reducing FLOPs by 13% compared to Bottleneck + attention (4.0G vs. 4.6G) and doubling the throughput (1362 imgs/s vs. 1305 imgs/s). Notably, while Bottleneck + attention (using EfficientViT’s [36] LiteMLA linear attention) slightly improves accuracy (82.3%), its computational cost is significantly higher, highlighting SS2D’s superior balance between performance and efficiency. Simpler designs like Bottleneck + GELU or Bottleneck + DWConv7×7 achieve marginal accuracy gains (82.1%) but fail to match SS2D’s global feature capture capability. These results underscore that SS2D, when combined with bottleneck structures, optimally leverages hardware-friendly operations (e.g., linear complexity scans) to enhance both accuracy and speed, aligning with the efficiency goals of modern lightweight architectures.
Table 8: Comparsion Results of different module in Bottleneck
Global mixer Params(M) FLOPs(G) Top-1(%) TP.(imgs/s)
no Bottleneck Mamba 21.4 3.3 81.9 2255
Bottleneck + GELU 23.5 3.6 82.1 1976
Bottleneck + DWConv7$\times$7 23.6 4.0 82.1 1880
Bottleneck + attention 26.4 4.6 82.3 1305
Bottleneck + SS2D 25.4 4.0 83.1 1362
More InceptionMamba Heatmap
The ablation study of the heatmap shows that InceptionMamba outperforms ConvNeXt [3] and Inceptionnext [4]. For the input images of a dog, InceptionMamba produces heatmaps that more comprehensively and accurately highlight object regions. It effectively highlights the dog’s limbs. This indicates that the InceptionMamba architecture, likely due to its unique Mamba-based temporal processing combined with Inception-like multi-scale feature fusion, improves global context modeling. The enhanced heatmap coherence implies that InceptionMamba captures richer spatial relationships, presenting it as a potential state-of-the-art lightweight architecture with qualitative superiority in visual feature attribution, which also complements previous ablation results regarding efficiency-accuracy trade-off.
Input
ConvNeXt
Inceptionnext
InceptionMamba
Figure 5: Comparison of heatmaps generated by different models.
5 Conclusion
In this paper, we introduce a novel hierarchical vision architecture named InceptionMamba through the lens of multi-scale spatial global interaction. Built upon the InceptionNeXt [4] framework, we present a modular architecture that leverages Large Band Depthwise Convolution as the Local Mixer for extracting local features and adopts Mamba [13, 14] Bottleneck as Global Mixer for global modeling, thereby enabling the network to capture local details and long-range semantic dependencies across hierarchical token spaces. Extensive experiments demonstrate the consistent superiority of InceptionMamba in terms of both accuracy and computational efficiency compared to representative ConvNets (e.g., ConvNeXt [3]) and Transformer (e.g., Swin [30]) on benchmarks spanning image classification and several downstream vision tasks. We hope our work can inspire the exploration of efficient multi-scale-global interaction paradigms in deep architecture design and facilitate the development of lightweight yet powerful vision backbones.